5,100 dollars invested for 10 years at 9 percent compounded annually is?
1
Expert's answer
2020-04-01T14:54:33-0400

We need to find the compound interest.

Given,


"P = \$5100 \\
t= 10 \space years \\
r=9\% = 0.09 \\
compounded \space Annually, so \space n = 1"

Formula to find the compound interest "= Amount - Principal"

"= P (1 + \frac{r}{n})^{nt} - P"



"= 5100 ( 1+ \frac{0.09}{1})^{10 \times 1} -5100 = 5100 (1.09)^{10} - 5100"

"= 5100 \times 2.3673 - 5100 = 12073.23 - 5100 = 6973.23\\"

"Amount=compound\, interest+Principal"

Answer:

The compounded interest = "\$ 6973.23"